# Any-Time Regret-Guaranteed Algorithm for Control of Linear Quadratic Systems

### v2

@@ 3 Any-time Regret SDP-based LQ cOntrol (ARSLO) algorithm
It can be shown that excessive updates in the policy can cause inefficiency, so there is a need for a criterion to decide on policy updates. By choosing $\det(V_{t})>(1+\beta)\det(V_{\tau})$ where $V_{\tau}$ is the covariance matrix of the most recently updated policy and arbitrarily setting $\beta=1$, we trigger a policy update when the ellipsoid volume is roughly halved. The parameters $\lambda$, $\Gamma_{t}$, and $\bar{\epsilon}(\bar{\kappa}_{1})$ are jointly tuned to ensure that the closed-loop system is $(\kappa_{*},\gamma_{*})$-strongly sequentially stable where $\kappa_{*}:=\sqrt{2\|P_{*}\|/\alpha_{0}}$ and $\gamma_{*}=\kappa_{*}^{-2}/2$, as demonstrated in Theorem 1. The rationale for choosing Gaussian perturbation noise with a covariance matrix $\Gamma_{t}$ such that $\|\Gamma_{t}\|=\mathcal{O}(1/\sqrt{t})$ is to ensure stability and achieve an overall regret of order $\mathcal{O}(\sqrt{t})$. Formally, the reasoning behind this choice is elaborated in Proposition 2. The algorithm enjoys any-time regret guarantees, as neither the initial estimate $\Theta_{0}$, the regularization parameter $\lambda$, nor the input perturbation noise $\eta_{t}$ depend on the time horizon of the algorithm’s execution.
@@ 3.1 Stability Analysis of ARSLO algorithm
     Compute $K(\mathcal{C}_{t})$ by (13) and through solving the relaxed primal SDP problem (12)
@@
$\eta_{t}\sim\mathcal{N}\big(0,\Gamma_{t}\big),\qquad\Gamma_{t}:=\frac{\bar{p}_{t}\sigma_{\omega}^{2}\Big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|}{\alpha_{0}}\Big)}{\sqrt{t+\bar{c}}},$ (22)
@@
As an immediate consequence of Theorem 1, the following corollary concludes the required tightness of the initial estimate $\Theta_{0}$.
@@
For the ARSLO algorithm to generate strongly sequentially stabilizing policies, the initial estimate $\Theta_{0}$ should lie within a $\bar{\epsilon}(\bar{\kappa}_{1})$-neighborhood of $\Theta_{*}$ of order $\mathcal{O}\!\left({1}/{\kappa_{*}^{14}}\right)$.
@@ 3.2 Regret Bound Analysis of ARSLO algorithm
where $\operatorname{Tr}(P_{*})\leq{D}$ is an a priori known bound. Unlike the LAG-LQ algorithm, our method does not require such an a priori bound; accordingly, no such quantity appears in the regret bound of ARSLO. This is indeed achieved by jointly leveraging the relaxed primal and dual SDP solutions when tuning the input perturbation covariance matrix, and by employing a carefully designed warm-up phase that repeatedly solves the relaxed dual SDP problem while using the computed upper bound for the operator norm of the DARE solution as given in (26). In this respect, our algorithm offers a clear advantage over the method of abeille2020efficient, in addition to its primary strength of coming with anytime regret guarantees. Moreover, the use of both primal and dual solutions further facilitates non-isotropic exploration. As for the dependence on the system ambient dimension and the operator norm of the DARE solution, note that, even in favor of the LAQ-LQ algorithm, if we let $D=\operatorname{Tr}(P_{*})\leq n\|P_{*}\|$, it has a dependence of $\mathcal{O}\!\bigl(\sqrt{n^{4}(n+m)^{2}\|P_{*}\|^{7}}\bigr)$. On the other hand, the dependence of the regret of the ARSLO algorithm on the dimension and on the solution of the DARE is $\mathcal{O}\bigl(\sqrt{n^{2}(n+m)^{3}\|P_{*}\|^{16}}\bigr)$. Leaving aside the fact that we conduct this comparison in favor of the LAG-LQ algorithm by setting $D=\operatorname{Tr}(P_{*})$, the results nevertheless indicate that the dependence on the dimension for ARSLO is of comparable order, or in some cases even better. However, our algorithm exhibits a stronger dependence on $\|P_{*}\|$, suggesting that there may still be room for improvement.
@@
The strong sequential-stability requirement mandates that each individually generated policy be stabilizing and that consecutively generated policies under the ARSLO algorithm remain sufficiently close, namely, $\|P^{-1/2}(\mathcal{C}_{t+1})P^{1/2}(\mathcal{C}_{t})\|\leq 1+\gamma_{*}/2$. Satisfying these condition necessitates a conservative algorithmic design, implemented by enforcing
$\mu_{t}\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{32{\kappa_{t}^{8}}}I$ (30)
@@
$\mu_{t}\,\|P(\mathcal{C}_{t})\|_{*}V_{t}^{-1}\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I.$ (31)
for any $t$ and some $\bar{\rho}\in[0,\,8)$. Enforcing this condition guarantees that any policy generated by the algorithm is $(\kappa_{*},\gamma_{*})$–strongly stabilizing and that consecutively generated policies remain within a controlled distance of one another. However, this distance requirement is weaker than that imposed by strong sequential stability, which corresponds to the case $\bar{\rho}=8$. Under this relaxation, the closed-loop system’s state norm may not remain adequately controlled and can exhibit undesired spikes under arbitrary switching. As established in switched-systems control theory, switching among individually stabilizing policies can still induce instability or lead to temporary explosions of the state norm. A standard remedy in this setting is to sufficiently slow down the switching. Motivated by this idea, for a given $\bar{\rho}\in[0,,8)$, the frequency of policy updates must be appropriately limited. This is achieved by tuning the exploration–exploitation trade-off parameter $\beta(\bar{\rho})$ in the update criterion
$\det(V_{t})>(1+\beta(\bar{\rho}))\det(V_{\tau}).$
Viewing $\bar{\rho}=8$, strong sequential stability as one extreme, the case $\bar{\rho}=0$ represents the opposite extreme, imposing only the minimal requirement that each generated policy be individually strongly stabilizing. The key question, then, is which value of $\bar{\rho}$ achieves the optimal regret. Intuitively, setting $\bar{\rho}$ to a smaller value increases the distance between consecutively generated policies $\|P^{-1/2}(\mathcal{C}_{t^{\prime}})P^{1/2}(\mathcal{C}_{t})\|\leq 1+2{\kappa_{*}^{3-\bar{\rho}/2}}$ (by Lemma 24 and Remark 4), which requires a longer dwell-time between policy updates, i.e., larger value of $\beta(\bar{\rho})$, potentially increasing the upper bound of the regret term $R_{4}(t)$. Conversely, a larger value of $\bar{\rho}$ increases the order of dependence of the perturbation-noise covariance matrix $\Gamma_{t}$ on $\|P_{*}\|$, which raises the upper bound of the regret term $R_{6}(t)$. This clearly illustrates the inherent trade-off and motivates a general analysis to identify the optimal value $\bar{\rho}_{*}$ that achieves the best regret bound in terms of its dependence on $\|P_{*}\|$. This analysis also allows for a careful examination of the trade-off between the regret bound and the high-probability state-norm guarantees. The rationale for specifying $\beta(\bar{\rho})$ is provided in (40) of Theorem 4.
@@
Fix $\bar{\rho}\in[0,\,8)$. Let Algorithm 3, set the regularization parameter $\lambda$ as
@@
$\eta_{t}\sim\mathcal{N}\big(0,\Gamma_{t}\big),\qquad\Gamma_{t}:=\frac{\bar{p}_{t}\sigma_{\omega}^{2}\Big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|}{\alpha_{0}}\Big)}{\sqrt{t+\bar{c}}},$ (36)
@@
[…]
@@
where $\kappa_{1}=\sqrt{2\|P(\mathcal{C}_{1})\|/\alpha_{0}}$. Then, under the above update mechanism, the closed-loop state is bounded as follows:
@@
where $\xi(\kappa_{*},\bar{\rho})=2{\kappa_{*}^{3-\bar{\rho}/2}}$.
@@ 5 Warm-up phase
$\|\Theta_{0}-\Theta_{*}\|_{F}\leq\epsilon.$
For ARSLO, it has been shown that $\epsilon=\bar{\epsilon}(\bar{\kappa}_{1})$, where $\bar{\epsilon}(\bar{\kappa}_{1})$ is given in (24), and
@@
For both cases, $\kappa_{1}=\sqrt{{2\|P(\mathcal{C}_{1})\|}/{\alpha_{0}}}$ where $P(\mathcal{C}_{1})$ is the solution of the relaxed dual SDP program associated with the initial ellipsoid $\mathcal{C}_{1}$. This ellipsoid is precisely the desired confidence ball $\|\Theta-\Theta_{0}\|_{F}\leq\epsilon$. Therefore, the warm-up phase must perform sufficient system identification to produce such a confidence ball and verify that this ball lies inside the required initial neighborhood. The following algorithm describes this warm-up procedure.
Algorithm 2, employs an initial strongly stabilizing policy $K_{0}$ and
@@
  Play input $u_{t}=K_{0}x_{t}+\nu_{t}$
@@
  Compute the least-squares estimate $\hat{\Theta}_{t}^{w}$ via (45)
  Compute the radius $r_{w}(t)$ via (46) and use it to compute $\epsilon_{w}(t)$ by (47)
  Construct the confidence ellipsoid (ball) $\mathcal{C}_{t}^{w}$ centered at $\hat{\Theta}_{t}^{w}$ with radius $\epsilon_{w}(t)$
@@
applies the input $u_{t}=K_{0}x_{t}+\nu_{t}$, where $\nu_{t}\sim\mathcal{N}(0,2\sigma_{\omega}^{2}\kappa_{0}^{2}I)$. It utilizes the saved data $z_{t}$ and $x_{t}$ to construct a confidence ellipsoid, similar to the one described in Section 2.2, with the difference that the least-squares estimation is regularized with respect to $\|\Theta\|^{2}_{F}$ using the regularization parameter $\rho$. The center of this ellipsoid is computed by
$\hat{\Theta}_{t}^{w}$ $=\bar{V}_{t}^{-1}\left(Z_{t}^{\top}X_{t}\right)$ (45)
$\bar{V}_{t}$ $=\rho I+Z_{t}^{\top}Z_{t}$
and
$r^{w}_{t}=\bigg(\sigma_{\omega}\sqrt{2n\log\frac{n\det(\bar{V}_{t})}{\delta\det(\lambda I)}}+\sqrt{\rho}\psi_{\Theta_{*}}\bigg)^{2}.$ (46)
where $Z_{t}$ and $X_{t}$ are constructed as described in the Section 2.2. Also, $\|\Theta_{*}\|_{F}\leq\psi_{\Theta_{*}}$, where one may set $\psi_{\Theta_{*}}=\sqrt{n}\vartheta$. The regularization parameter is chosen as $\rho=\sigma_{\omega}^{2}\vartheta^{-2}$.
@@
Theorem 6
Let $K_{0}$ be a $(\kappa_{0},\gamma_{0})$-strongly stabilizing policy and let $\nu_{t}\sim\mathcal{N}(0,2\sigma_{\omega}^{2}\kappa_{0}^{2}I)$ be an input perturbation. Then, Algorithm 2 (the ‘warm-up phase’), by applying the control input $u_{t}=K_{0}x_{t}+\nu_{t}$, outputs $\Theta_{0}$ such that $\|\Theta_{0}-\Theta_{*}\|_{F}\leq\varepsilon$ if the following condition met
$\epsilon_{w}(t)=\sqrt{\frac{r^{w}_{t}}{\lambda_{min}(V_{t})}}\leq\varepsilon$ (47)
where
[…]
The following remark specifies the scaling of the warm-up duration $t_{w}$.

### v3

@@ 3 Any-time Regret SDP-based LQ cOntrol (ARSLO) Algorithm
It can be shown that excessive updates in the policy can cause inefficiency, so there is a need for a criterion to decide on policy updates. By choosing $\det(V_{t})>(1+\beta)\det(V_{\tau})$ where $V_{\tau}$ is the covariance matrix of the most recently updated policy and arbitrarily setting $\beta=1$, we trigger a policy update when the ellipsoid volume is roughly halved. The parameters $\lambda$, $\Gamma_{t}$, and $\bar{\epsilon}(\bar{\kappa}_{1})$ are jointly tuned to ensure that the closed-loop system is $(\kappa_{*},\gamma_{*})$-strongly sequentially stable where $\kappa_{*}:=\sqrt{2\|P_{*}\|/\alpha_{0}}$ and $\gamma_{*}=\kappa_{*}^{-2}/2$, as demonstrated in Theorem 1. The rationale for choosing Gaussian perturbation noise with a covariance matrix $\Gamma_{t}$ such that $\|\Gamma_{t}\|=\mathcal{O}(1/\sqrt{t})$ is to ensure stability and achieve an overall regret of order $\mathcal{O}(\sqrt{t})$. Formally, the reasoning behind this choice is elaborated in Appendix M. The algorithm enjoys any-time regret guarantees, as neither the initial estimate $\Theta_{0}$, the regularization parameter $\lambda$, nor the input perturbation noise $\eta_{t}$ depend on the time horizon of the algorithm’s execution.
@@ 3.1 Stability Analysis of ARSLO Algorithm
     Solve the relaxed primal SDP problem (12) and compute $K(\mathcal{C}_{t})$ by (13)
@@
$\eta_{t}\sim\mathcal{N}\big(0,\Gamma_{t}\big),\qquad\Gamma_{t}:=\frac{\bar{p}_{t}\sigma_{\omega}^{2}\Big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|}{\alpha_{0}}I\Big)}{\sqrt{t+\bar{c}}},$ (22)
@@
As an immediate consequence of Theorem 1, and in particular the requirement (24), the following corollary characterizes the order of magnitude of $\bar{\epsilon}(\bar{\kappa}_{1})$.
@@
For the ARSLO algorithm to generate strongly sequentially stabilizing policies, the initial estimate $\Theta_{0}$ must lie within a $\bar{\epsilon}(\bar{\kappa}_{1})$-neighborhood of $\Theta_{*}$, with $\bar{\epsilon}(\bar{\kappa}_{1})=\mathcal{O}\!\left(1/\kappa_{*}^{14}\right)$.
@@ 3.2 Regret Bound Analysis of ARSLO Algorithm
where $\operatorname{Tr}(P_{*})\leq{D}$ is an a priori known bound. Unlike the LAG-LQ algorithm, our method does not require such an a priori bound; accordingly, no such quantity appears in the regret bound of ARSLO. This is indeed achieved by jointly leveraging the relaxed primal and dual SDP solutions when tuning the input perturbation covariance matrix, and by employing a carefully designed warm-up phase that repeatedly solves the relaxed dual SDP problem while using the computed upper bound for the operator norm of the DARE solution as given in (26). In this respect, our algorithm offers a complementary viewpoint to the method of abeille2020efficient, in addition to its primary strength of coming with anytime regret guarantees. Moreover, the use of both primal and dual solutions further facilitates non-isotropic exploration. As for the dependence on the system ambient dimension and the operator norm of the DARE solution, note that, even in favor of the LAQ-LQ algorithm, if we let $D=\operatorname{Tr}(P_{*})\leq n\|P_{*}\|$, it has a dependence of $\mathcal{O}\!\bigl(\sqrt{n^{4}(n+m)^{2}\|P_{*}\|^{7}}\bigr)$. On the other hand, the dependence of the regret of the ARSLO algorithm on the dimension and on the solution of the DARE is $\mathcal{O}\bigl(\sqrt{n^{2}(n+m)^{3}\|P_{*}\|^{16}}\bigr)$. Leaving aside the fact that we conduct this comparison in favor of the LAG-LQ algorithm by setting $D=\operatorname{Tr}(P_{*})$, the results nevertheless indicate that the dependence on the dimension for ARSLO is of comparable order, or in some cases even better. However, our algorithm exhibits a stronger dependence on $\|P_{*}\|$, suggesting that there may still be room for improvement.
@@
The strong sequential-stability requirement mandates that each individually generated policy be stabilizing and that consecutively generated policies under the ARSLO algorithm remain sufficiently close, namely, $\|P^{-1/2}(\mathcal{C}_{t+1})P^{1/2}(\mathcal{C}_{t})\|\leq 1+\gamma_{*}/2$. Satisfying this condition necessitates a conservative algorithmic design, implemented by enforcing
$\mu_{t}\|P(\mathcal{C}_{t})\|V_{t}^{-1}\preceq\frac{\alpha_{0}}{32{\kappa_{t}^{8}}}I$ (30)
@@
$\mu_{t}\,\|P(\mathcal{C}_{t})\|V_{t}^{-1}\preceq\frac{\alpha_{0}}{4\kappa_{t}^{\bar{\rho}}}I.$ (31)
for any $t$ and some $\bar{\rho}\in[0,\,8)$. Enforcing this condition guarantees that any policy generated by the algorithm is $(\kappa_{*},\gamma_{*})$–strongly stabilizing and that consecutively generated policies remain within a controlled distance of one another. However, this distance requirement is weaker than that imposed by strong sequential stability, which corresponds to the case $\bar{\rho}=8$. Under this relaxation, the closed-loop system’s state norm may not remain adequately controlled and can exhibit undesired spikes under arbitrary switching. As established in switched-systems control theory, switching among individually stabilizing policies can still induce instability or lead to temporary explosions of the state norm. A standard remedy in this setting is to sufficiently slow down the switching. Motivated by this idea, for a given $\bar{\rho}\in[0,\,8)$, the frequency of policy updates must be appropriately limited. This is achieved by tuning the exploration–exploitation trade-off parameter $\beta(\bar{\rho})$ in the update criterion
$\det(V_{t})>(1+\beta(\bar{\rho}))\det(V_{\tau})$
with $\bar{\rho}=8$ corresponding to strong sequential stability while the other extreme, $\bar{\rho}=0$, imposes the minimal requirement that each generated policy be individually strongly stabilizing. The key question, then, is which value of $\bar{\rho}$ achieves the optimal regret. Intuitively, setting $\bar{\rho}$ to a smaller value increases the distance between consecutively generated policies $\|P^{-1/2}(\mathcal{C}_{t^{\prime}})P^{1/2}(\mathcal{C}_{t})\|\leq 1+2{\kappa_{*}^{3-\bar{\rho}/2}}$ (by Lemma 24 and Remark 4), which requires a longer dwell-time between policy updates, i.e., larger value of $\beta(\bar{\rho})$, potentially increasing the upper bound of the regret term $R_{4}(t)$. Conversely, a larger value of $\bar{\rho}$ increases the order of dependence of the perturbation-noise covariance matrix $\Gamma_{t}$ on $\|P_{*}\|$, which raises the upper bound of the regret term $R_{6}(t)$. This clearly illustrates the inherent trade-off and motivates a general analysis to identify the optimal value $\bar{\rho}_{*}$ that achieves the best regret bound in terms of its dependence on $\|P_{*}\|$. This analysis also allows for a careful examination of the trade-off between the regret bound and the high-probability state-norm guarantees. The rationale for specifying $\beta(\bar{\rho})$ is provided in (40) of Theorem 4.
@@
Fix $\bar{\rho}\in[0,\,8)$. Set the regularization parameter $\lambda$ in Algorithm 3 to
@@
$\eta_{t}\sim\mathcal{N}\big(0,\Gamma_{t}\big),\qquad\Gamma_{t}:=\frac{\bar{p}_{t}\sigma_{\omega}^{2}\Big(K(\mathcal{C}_{t})K^{\top}(\mathcal{C}_{t})+\frac{\|P(\mathcal{C}_{t})\|}{\alpha_{0}}I\Big)}{\sqrt{t+\bar{c}}},$ (36)
@@
Let $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$, $\lambda$, and $\Gamma_{t}$ be tuned according to Theorem 3. Suppose the algorithm updates the policy whenever the criterion $\det(V_{t})>(1+\beta(\bar{\rho}))\det(V_{\tau})$” is triggered, with $\tau$ being the last time of policy update. Let
@@
where $\kappa_{1}=\sqrt{2\|P(\mathcal{C}_{1})\|/\alpha_{0}}$. Then, the closed-loop state satisfies
@@
with probability at least $1-\delta$ where $\xi(\kappa_{*},\bar{\rho})=2{\kappa_{*}^{3-\bar{\rho}/2}}$.
@@
As an immediate consequence of Theorem 3, and in particular the requirement (38), the following corollary characterizes the order of magnitude of $\underline{\epsilon}(\bar{\kappa}_{1},\bar{\rho})$.
@@ 5 Warm-up Phase
$\|\Theta_{0}-\Theta_{*}\|_{F}\leq\varepsilon.$
For ARSLO, it has been shown that $\varepsilon=\bar{\epsilon}(\bar{\kappa}_{1})$, where $\bar{\epsilon}(\bar{\kappa}_{1})$ is given in (24), and
@@
  Play input $u_{t}=K_{0}x_{t}+\nu_{t}$ where $\nu_{t}\sim\mathcal{N}(0,2\sigma_{\omega}^{2}\kappa_{0}^{2}I)$
@@
  Compute $\bar{V}_{t}$ and then $\hat{\Theta}_{t}^{w}$ using (45), followed by $r_{t}^{w}$ using (47)
  Compute $\epsilon_{w}(t)$ using (49) and construct the confidence ball $\mathcal{C}_{t}^{w}$
@@
For both cases, $\kappa_{1}=\sqrt{{2\|P(\mathcal{C}_{1})\|}/{\alpha_{0}}}$ where $P(\mathcal{C}_{1})$ is the solution of the relaxed dual SDP program associated with the initial ellipsoid (ball) $\mathcal{C}_{1}$. This ellipsoid is precisely the desired initial estimate, $\|\Theta-\Theta_{0}\|_{F}\leq\varepsilon$. Therefore, the warm-up phase must perform sufficient system identification to produce an initial estimate $\Theta_{0}$ and verify that it lies inside the required initial neighborhood.
Algorithm 2, employs an initial $(\kappa_{0},\gamma_{0})$-strongly stabilizing policy $K_{0}$ and applies the input $u_{t}=K_{0}x_{t}+\nu_{t}$, where $\nu_{t}\sim\mathcal{N}(0,2\sigma_{\omega}^{2}\kappa_{0}^{2}I)$. Using the collected data pairs $\{(z_{t},x_{t})\}$, the algorithm constructs a confidence ellipsoid of the form
$\operatorname{Tr}\!\left((\Theta-\hat{\Theta}^{w}_{t})^{\top}\bar{V}_{t}(\Theta-\hat{\Theta}^{w}_{t})\right)\leq r_{t}^{w},$
which is analogous to the ellipsoid described in Section 2.2, except that the least-squares estimator is regularized with respect to $\|\Theta\|_{F}^{2}$ using a regularization parameter $\rho$. The center of the confidence ellipsoid is computed by
$\hat{\Theta}_{t}^{w}$ $=\bar{V}_{t}^{-1}Z_{t}^{\top}X_{t},$ (45)
$\bar{V}_{t}$ $=\rho I+Z_{t}^{\top}Z_{t},$ (46)
where $Z_{t}$ and $X_{t}$ are constructed as described in Section 2.2. The radius parameter $r_{t}^{w}$ is defined as
$r_{t}^{w}=\left(\sigma_{\omega}\sqrt{2n\log\frac{n\det(\bar{V}_{t})}{\delta\det(\rho I)}}+\sqrt{\rho}\,\psi_{\Theta_{*}}\right)^{2},$ (47)
where in construction we applied $\|\Theta_{*}\|_{F}\leq\psi_{\Theta_{*}}$. In particular, one may set $\psi_{\Theta_{*}}=\sqrt{n}\,\vartheta$, in which case the regularization parameter is chosen as $\rho=\sigma_{\omega}^{2}\vartheta^{-2}$.
The largest confidence ball contained in this ellipsoid is then defined as
$\mathcal{C}_{t}^{w}(\delta):=\left\{\Theta\,\middle|\,\|\Theta-\hat{\Theta}_{t}^{w}\|_{F}\leq\epsilon_{w}(t)\right\},$ (48)
where $\Theta_{*}$ belongs to it with probability at least $1-\delta$ with
$\epsilon_{w}(t)=\sqrt{\frac{r_{t}^{w}}{\lambda_{\min}(\bar{V}_{t})}}.$ (49)